OpenStudy (anonymous):

ALGEBRA HELP PLEASEEE Find the GCF of the terms of each polynomial. Then factor the polynomial. 22. 12p^4 + 16p^3 + 8p = 24. 45c^5 - 63c^3 + 27c =

OpenStudy (anonymous):

im guessing its taking out the common terms then simplifying

OpenStudy (jdoe0001):

\(12p^4 + 16p^3 + 8p\qquad \begin{cases} 12=3\cdot 2\cdot 2\\ 16=2\cdot 2\cdot 2\cdot 2\\ 8=2\cdot 2\cdot 2\\ p^4=pppp\\ p^3=ppp \end{cases}\impliedby hint\)

OpenStudy (jdoe0001):

\(45c^5 - 63c^3 + 27c \qquad \begin{cases} 45=3\cdot 3\cdot 5\\ 63=7\cdot 3\cdot 3\\ 27=3\cdot 3\cdot 3\\ c^5=ccccc\\ c^3=ccc \end{cases}\impliedby hint\)

OpenStudy (anonymous):

4p • (3p^3 + 4p^2 + 2) is this correct?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`4p • (3p^3 + 4p^2 + 2) is this correct?` yep, 12p^4 + 16p^3 + 8p = 4p(3p^3 + 4p^2 + 2)

OpenStudy (anonymous):

@jim_thompson5910 Okay and 4p • (3p^3 + 4p^2 + 2) is in standard form?

jimthompson5910 (jim_thompson5910):

it's in factored form

OpenStudy (anonymous):

Sorry, that's what I meant. Thank you

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

Can you help me with a few more please?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

For #24 for factored form I got this: 9c • (5c^4- 7c^2 + 3)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Okay, thank you. Simplify each product. Write in standard form. 38. (y - 6) (3y + 7) = 40. (4b - 3) (4b + 3) =

jimthompson5910 (jim_thompson5910):

for those, I would use the box method